Using the method of undetermined coefficients, solve the equation:


d²y/dx² -3dy/dx +2y=4x²

1
Expert's answer
2022-01-24T03:14:23-0500

"\frac{d²y}{dx²} - 3 \frac{dy}{dx} + 2y=4x²\\


\Big\{\frac{d²}{dx²} - 3 \frac{d}{dx} + 2\Big\}y=4x²\\

\text{Let} \, m = \frac{d}{dx} \\

\text{Then, we have the auxiliary equation}\\

m^2 - 3m + 2 = 0 \\

m^2 - m - 2m + 2 = 0 \\

m(m - 1) - 2(m - 1) = 0\\

(m - 2)(m - 1) = 0\\

m = 2,1 \\

\text{Hence, we have the complementary function}\\

C.F = C_1 e^x + C_2 e^{2x} \\
\text{where} \, C_1 \text{and} \, C_2 \, \text{are arbitrary constants} \\

\text{To find the particular integeral by}\\
\text{the method of undetermined coefficient}\\

\text{Let the particular integral be of the form}\\

y_p = ax^2 + bx + c \\

\text{We want to find a, b, c such that}\\
y''_p - 3y'_p + 2 y_p = 4x^2 . . . (*) \\

\text{From} \, y_p = ax^2 + bx + c \, \text{.We have}\\

y'p = 2ax + b \\

y''p = 2a \\

\text{Now, substitute these values into eq(*)}\\
\text{ To have,}\\

2a - 3(2ax + b) + 2(ax^2 + bx + c) = 4x^2 \\

\text{Comparing powers. We have}\\

2ax^2 = 4x^2 \, \implies \, 2a = 4 \, \implies a = 2\\

(-6a + 2b)x = 0x \, \implies -12 + 2b = 0\\ \implies b = 6\\

2a - 3b + 2c = 0 \, \implies 4 - 18 + 2c = 0\\ \implies 2c = 14 \implies c = 7\\

\text{Hence, }\\

\text{P.I}\, = y_p = ax^2 + bx + c = 2x^2 + 6x + 7\\

\text{Thus, the general solution is}\\

y(x) = C.F + P.I \\

y(x) = C_1 e^x + C_2 e^{2x} + 2x^2 + 6x + 7 \\
\text{where} \, C_1 \, \text{and} \, C_2 \, \text{are arbitrary constants}"